OpenStudy (anonymous):

how do you solve x/4 = 2 + (x-3)/3 (the slashes mean over)

OpenStudy (anonymous):

3x/4=2+(x-3) 3x=8+4x-12 3x-4x=-4 -x=-4 x=4